# Structure-reinforced Transformer for Dynamic Graph Representation Learning with Edge Temporal States

### v2

@@ Abstract
The burgeoning field of dynamic graph representation learning, fuelled by the increasing demand for graph data analysis in real-world applications, poses both enticing opportunities and formidable challenges. Despite the promising results achieved by recent research leveraging recurrent neural networks (RNNs) and graph neural networks (GNNs), these approaches often fail to adequately consider the impact of the edge temporal states on the strength of inter-node relationships across different time slices, further overlooking the dynamic changes in node features induced by fluctuations in relationship strength. Furthermore, the extraction of global structural features is hindered by the inherent over-smoothing drawback of GNNs, which in turn limits their overall performance. In this paper, we introduce a novel dynamic graph representation learning framework namely Recurrent Structure-reinforced Graph Transformer (RSGT), which initially models the temporal status of edges explicitly by utilizing different edge types and weights based on the differences between any two consecutive snapshots. In this manner, the varying edge temporal states are mapped as a part of the topological structure of the graph. Subsequently, a structure-reinforced graph transformer is proposed to capture temporal node representations that encoding both the graph topological structure and evolving dynamics,through a recurrent learning paradigm. Our experimental evaluations, conducted on four real-world datasets, underscore the superior performance of the RSGT in the realm of discrete dynamic graph representation learning. The results reveal that RSGT consistently surpasses competing methods in dynamic link prediction tasks.
Index Terms:
@@
Dynamic Graph, Graph Representation Learning, Structure-reinforced Graph Transformer, Edge Temporal States.
I Introduction
@@
Graphs find extensive use across myriad application domains, such as traffic flow forecasting [1, 2], recommender systems [3, 4, 5] and stock prediction [6], to delineate the interactions between entities within real-world complex systems. As these systems inherently exhibit temporal variations, with entities and interactions frequently emerging or disappearing, they can be intuitively abstracted as either discrete [7, 8] or continuous dynamic graphs [9, 10]. Learning the representation of these dynamic graphs is of paramount importance as it encapsulates the interaction and evolution mechanisms of the graph, thereby aiding in understanding and predicting the system behavior. Contrasting with static graph representation [11, 12], the task of dynamic graph representation learning not only necessitates preserving the topological features but also reflecting the evolving dynamic characteristics of the graph. This requirement imposes significant complexity on the model design, rendering the task particularly challenging. In this study, we primarily concentrate on the discrete dynamic graph representation learning task, owing to its prevalence in various applications [7, 13, 8].
A discrete dynamic graph is modeled as a sequence of ordered static snapshots across discrete time intervals, capturing changes in graph topology, including node and edge modifications. Recent advancements [7, 13, 14] in discrete dynamic graph representation learning have focused on integrating Graph Neural Networks (GNNs) [15, 16] with sequence modeling techniques [17], which aims to leverage both the structural attributes of graphs and their temporal evolution. Typically, these approaches utilize GNNs to extract structural features from each time slice of the graph, while sequence models like LSTM and variational auto-encoders capture the dynamics of these structural features over time, updating them accordingly. Despite their effectiveness, these models have encountered challenges in inductive learning, particularly in incorporating new nodes based on temporal dynamics. To address these challenges, novel strategies have been proposed. These include evolving GNN parameters via RNNs [7], embedding snapshots using deep autoencoders [18], and employing temporal self-attention mechanisms [13]. Such approaches aim to enhance the simultaneous learning of graph structure and temporal dynamics, thereby improving the models’ inductive learning capabilities for dynamic graphs.
However, the existing approaches have several drawbacks that limit their effectiveness in downstream graph-oriented tasks. Firstly, these methods do not explicitly account for the dynamic influences of edge temporal states on node features. It is intuitively that a node’s characteristics are significantly shaped by its interactions with other neighbors that may persist, appear, or disappear over time across different time slices. These changes in edge temporal states and the varying strengths of relationships among nodes play a crucial role in the accurate learning of node representations [19, 20]. Yet, current research often neglects the critical impact of these variations of edge temporal states, overlooking a fundamental aspect of effective graph representation learning. Secondly, GNNs face challenges of over-smoothing [21, 22], which hinders the ability of GNNs to capture and differentiate between local and global structural features effectively, further constraining the quality of node representation learning. Overcoming these limitations requires innovative strategies that preserve the distinctiveness of node features while incorporating the dynamics of edge temporal states.
In this paper, we introduce a novel framework designed for the dynamic graph representation learning, named Recurrent Structure-reinforced Graph Transformer (RSGT). It innovatively transforms each graph snapshot into a weighted multi-relation graph, which not only highlights the edge temporal states but also delineates the variations between consecutive snapshots through distinct edge types and weights. Acknowledging these modifications in node features, stemming from snapshot differences, is vital for decoding the evolutionary patterns inherent in dynamic graphs, we explicitly account for such edge temporal states in our approach. We then conceptualize dynamic graph representation learning as a sequence of learning tasks associated with static graphs, implemented recurrently. Beginning with node features extracted from an initial snapshot, these features undergo continual updates in response to changes captured in subsequent snapshots. Drawing on the transformative impact of the Transformer model in static graph-related tasks [23, 24], we introduce a structure-reinforced graph transformer, which is adept at extracting both the nuanced local-global topological characteristics and the dynamic features evolving within the weighted multi-relation dynamic graph.
@@
Our extensive empirical evaluations, conducted across four real-world datasets, clearly demonstrate RSGT’s exceptional capabilities in the domain of discrete dynamic graph representation learning. Particularly in dynamic link prediction tasks, RSGT consistently outshines existing methods. These findings not only underscore RSGT’s superior performance but also highlight its potential to further dynamic graph analysis. In conclusion, the key contributions of this paper can be summarized as follows:
  • We propose a novel recurrent learning framework named RSGT, designed specially for dynamic graphs. It formulates a dynamic weighted multi-relational difference graph, which explicitly models the edge temporal states by different edge types and their associated weights across multiple time slices. By this way, we ingeniously integrate such temporal states into the graph topological structure, thereby seamlessly incorporating the edge temporal states into the dynamic graph representation learning.
  • A unique structure-reinforced graph transformer is designed for better node feature extraction. It is capable of concurrently extracting local-global topological features and evolving dynamic characteristics from the weighted multi-relation graph. This concurrent feature extraction significantly enhances the efficacy of the dynamic graph representation learning.
  • We have conducted comprehensive experiments on four real-world datasets to validate the performance of our RSGT framework. These experiments conclusively demonstrate its superior performance in comparison to existing methods.
The remainder of this paper is structured as follows. In Section II, we provide an essential preliminaries and formulates our research problem of dynamic graph representation learning. The approach proposed for the dynamic graph representation learning is comprehensively elaborated in Section III. Experimental results are presented and thoroughly analyzed in Section IV. We review the related work in Section V. Section VI draws conclusion and points out the future work.
II Problem Formulation
@@
Here, we present the formulations for dynamic graph representation learning through a set of definitions. A brief summary of the frequently used notations can be found in Table I.
TABLE I: Frequently used notations in this paper.
Notations Description
$\mathcal{G}$ A dynamic graph that changes discretely over time
$\mathcal{G}_{t}$ A specific snapshot of the dynamic graph $\mathcal{G}$ at a discrete time slice $t$
$V_{t}$ The set of vertices present in the snapshot $\mathcal{G}_{t}$
$|V_{t}|$ The cardinality of set $V_{t}$, indicating the number of vertices in snapshot $\mathcal{G}_{t}$
$E_{t}$ The set of edges present in the snapshot $\mathcal{G}_{t}$
$W_{t}$ The set of weights associated with the edges in the snapshot $\mathcal{G}_{t}$
$\textbf{H}^{t}$ The matrix representation of $\mathcal{G}_{t}$ in a lower-dimensional space
$d$ The number of feature dimensions in the target embedding space
X The initial matrix that embeds vertices from the dynamic graph $\mathcal{G}$
$\textbf{I}_{t-1}$ The latent temporal state derived from previous snapshots $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t-1}\}$
$f(\cdot|\Theta_{f})$ A neural network that is parameterized by a set of trainable parameters $\Theta_{f}$
@@ X The initial matrix that embeds vertices from the dynamic graph $\mathcal{G}$
A discrete dynamic graph, denoted as $\mathcal{G}$, can be conceptualized as an ordered sequence of static graphs represented as $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t}\}$. In this representation, $\mathcal{G}_{t}$ corresponds to a snapshot of the dynamic graph $\mathcal{G}$ at a specific time slice $t$. Each of these static graphs, $\mathcal{G}_{t}$, can be described as a triplet $<V_{t},E_{t},W_{t}>$, where $V_{t}$ signifies the set of nodes at time $t$, $E_{t}$ represents the set of edges, and $W_{t}$ corresponds to the set of edge weights. If an edge $e_{ij}^{t}$ exists in $E_{t}$, it implies that there is a link between node $v_{i}\in V_{t}$ and node $v_{j}\in V_{t}$ at time slice $t$, with an associated edge weight of $w_{ij}^{t}\in W_{t}$.
The temporal evolution of a graph, characterized by the appearance or disappearance of nodes and edges, necessitates the use of dynamic graph representation learning. This task aims to understand the evolutionary patterns of the graph to better capture node representations.
@@
Dynamic graph representation learning is a task that seeks to devise a mapping function capable of deriving a low-dimensional representation $\textbf{H}^{t}\in\Re^{|V_{t}|\times d}$ ($d\ll|V_{t}|$), which is achieved by utilizing a succession of snapshots $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t}\}$ until the time slice $t$. The $i$-th row of $\textbf{H}^{t}$, represented as $\textbf{h}^{t}_{i}$, signifies the derived representation of node $v_{i}$ at time slice $t$, encapsulating both the structural information and evolutional dynamics.
A typical learning framework can be articulated as follows:
$\textbf{H}^{t}=g(f(\mathcal{G_{t}},\textbf{X}|\Theta_{f}),\textbf{I}_{t-1}|\Theta_{g})$ (1)
where X denotes the initial node features or embedding matrix, $\textbf{I}_{t-1}$ symbolizes the most recent latent temporal state, $f$ designates an encoding model tasked with extracting structural features, and $g$ represents a sequence model responsible for discerning evolving patterns.
The effectiveness of node representations plays a pivotal role in tasks associated with graph orientation, such as dynamic link prediction [25, 10], node classification [26], etc. In this study, we utilize dynamic link prediction as an auxiliary task to bolster the model training of our proposed RSGT, and subsequently evaluate the effectiveness of the node representations obtained.
@@
Given a series of snapshots $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t}\}$ till time slice $t$, dynamic link prediction task is to predict whether a candidate edge $e_{ij}^{t+1}$ exists in the upcoming snapshot $\mathcal{G}_{t+1}$ at time slice $t+1$. which can be formulated as:
$p_{ij}^{t+1}=\tau(\mathcal{G}_{t+1}|\mathcal{G}_{1},\dots,\mathcal{G}_{t};\Theta_{\tau})$ (2)
@@
where $p_{ij}^{t+1}$ represents the predicted probability that $e_{ij}\in E_{t+1}$, $\tau$ indicates the prediction model.
III Approach
Figure 1: Overall framework of the recurrent structure-reinforced graph transformer.
@@ III Approach
The overall framework of RSGT is depicted in Figure 1. The framework primarily consists of two stages: (a) edge temporal state modeling for dynamic graph, which is focused on the assignment of diverse relational types and weights to edges. This assignment is based on the variabilities observed between adjacent snapshots; (b) recurrent temporal node feature extraction, in which a structure-reinforced graph transformer is employed. This transformer takes into account both the graph topology and the edge temporal states over the prevailing weighted multi-relation graph. The purpose of this is to recurrently update the node features. In the subsequent sections, we delve deeper into the intricacies of each stage.
Figure 2: Structure-reinforced graph transformer for temporal node feature extraction.
III-A Edge Temporal State Modeling of Dynamic Graph
@@
In a dynamic graph $\mathcal{G}$, it is intuitive to perceive the feature of a focal node $v_{i}$ as a reflection of its interactions with other nodes $v_{j}\in(N_{i}^{h})^{t}$ that it directly or indirectly engages with [27]. Here, $(N_{i}^{h})^{t}$ signifies the $h$-hop neighborhood of $v_{i}$ at a specific time slice $t$.
@@
It is observed from [19, 20] that social network theory suggests that the strength of the relationship between nodes predominantly depends on factors such as historical interaction frequency and duration. This relationship strength, in turn, shapes future interactions, thus playing a significant role in the formation and evolution of the graph. Consequently, capturing these dynamic interactions is instrumental in comprehending the dynamism of the graph. Observationally, the interactions between nodes exhibit persistence, emergence, or disappearance across multiple time slices, resulting in perpetual alterations in the temporal states of the edges. Therefore, understanding the evolution of $\mathcal{G}$ hinges on accurately learning how distinct edge temporal states impact node features across different time slice snapshots.
By explicitly modeling these edge temporal states, we can provide heuristic information to the downstream feature extraction model, thereby enabling effective learning of node representations. The temporal state of a specific edge is primarily characterized by two features: short-term dynamics, which represents the edge’s immediate persistence, emergence, or disappearance from the preceding snapshot, and long-term dynamics, which constitutes the aggregate impact of this edge on the strength of the correlation between its associated nodes.
In this paper, we propose an edge temporal state modeling strategy based on the weighted multi-relation difference graph. This strategy translates each snapshot, denoted as $\mathcal{G}_{t}$ at a given time slice $t$, into a weighted multi-relation graph, represented as $\hat{\mathcal{G}_{t}}=<V_{t},\hat{E}_{t},TP_{t},\hat{W}_{t}>$, where $TP_{t}$ symbolizes the set of edge types and $\hat{W}_{t}$ signifies the set of edge weights. The diversity of edge types encapsulates the short-term dynamics, while the associated edge weights capture the long-term dynamics.
@@
From eq. 3 and eq. 4, it can be observed that in $\hat{\mathcal{G}}_{t}$, we provisionally reconstruct the edges that vanished between time slices $t-1$ and $t$. This constitutes an intervention predicated on the premise that the disappearance of an edge does not necessarily signal an immediate termination of inter-influence between the connected nodes. Conversely, the disappearance of an edge might potentially provoke a spectrum of impacts on the two nodes. Consequently, it is crucial for subsequent feature extraction models to explicitly recognize this aspect, as it significantly contributes to comprehending the dynamic nature of the graph.
In examining the association between the duration of an edge and the correlation strength of nodes, edge weight temporal updating assigns a strength coefficient, denoted as $\omega$, to each edge. This coefficient is directly proportional to the duration of the edge, signifying that a longer edge duration corresponds to a higher strength coefficient:
@@
where $\omega^{t}_{ij}\in\hat{W}_{t}$ denotes the strength coefficient of the edge $e_{ij}^{t}$ at a specific time slice $t$, $k$ is indicative of the edge duration up to the point $t$. The terms $\alpha$ and $\beta$ are employed to denote the scale and power factors, respectively. As a result, the dynamic graph $\hat{\mathcal{G}_{t}}$ conserves the original topology while concurrently modeling the temporal state of the edge in a collaborative manner, taking into account the varying edge types and weights. This approach ensures a comprehensive representation of the edge dynamics across different time slices.
III-B Recurrent Temporal Node Feature Extraction
Figure 1 highlights the process of recurrent temporal node feature extraction over the constructed weighted multi-relation graph $\hat{\mathcal{G}}_{t}$ in its right section. The structure-reinforced graph transformer (SGT), a temporal node feature extraction model, takes the node feature $\textbf{H}^{t-1}\in\Re^{|V|\times d}$ from the previous time slice $t-1$ as its input. Here, $|V|$ signifies the maximum node count across all time slices, and $d$ represents the dimensionality of the temporal node representation.
The SGT updates node representations by integrating the structural information of $\hat{\mathcal{G}}_{t}$. This information amalgamates the graph topology with temporal edge states. Subsequently, $\textbf{H}^{t-1}$ is residually incorporated into the SGT’s output, $\hat{\textbf{H}}^{t}$, resulting in the updated node features $\textbf{H}^{t}\in\Re^{|V|\times d}$ at time slice $t$. The formulation of this procedure is as follows:
$\hat{\textbf{H}}^{t}=\mathcal{F}_{g_{l}}(\hat{\mathcal{G}}_{t},\textbf{H}^{t-1})$ (6)

### v3

# Dynamic Graph Representation Learning via Edge Temporal States Modeling and Structure-reinforced Transformer

@@ Abstract
Dynamic graph representation learning has emerged as a crucial research area, driven by the growing need for analyzing time-evolving graph data in real-world applications. While recent approaches leveraging recurrent neural networks (RNNs) and graph neural networks (GNNs) have shown promise, they often fail to adequately capture the impact of edge temporal states on inter-node relationships, consequently overlooking the dynamic changes in node features induced by these evolving relationships. Furthermore, these methods suffer from GNNs’ inherent over-smoothing problem, which hinders the extraction of global structural features. To address these challenges, we introduce the Recurrent Structure-reinforced Graph Transformer (RSGT), a novel framework for dynamic graph representation learning. It first designs a heuristic method to explicitly model edge temporal states by employing different edge types and weights based on the differences between consecutive snapshots, thereby integrating varying edge temporal states into the graph’s topological structure. We then propose a structure-reinforced graph transformer that captures temporal node representations encoding both graph topology and evolving dynamics through a recurrent learning paradigm, enabling the extraction of both local and global structural features. Comprehensive experiments on four real-world datasets demonstrate RSGT’s superior performance in discrete dynamic graph representation learning, consistently outperforming existing methods in dynamic link prediction tasks.
keywords:
@@
Dynamic Graph , Graph Representation Learning , Edge Temporal States , Structure-reinforced Graph Transformer
[…]
organization=School of Computer Engineering and Science, Shanghai University,city=Shanghai, postcode=200444, country=China
[…]
organization=School of Computer Science and Technology, Donghua University,city=Shanghai, postcode=201620, country=China
[…]
organization=School of Computer and Information Engineering, Shanghai Polytechnic University,city=Shanghai, postcode=201209, country=China
1 Introduction
@@
Graphs find extensive use across myriad application domains, such as traffic flow forecasting [1, 2], recommender systems [3, 4, 5] and stock prediction [6], to delineate the interactions between entities within real-world complex systems. As these systems inherently exhibit temporal variations, with entities and interactions frequently emerging or disappearing, they can be intuitively abstracted as dynamic graphs. Dynamic graphs, which capture the evolution of relationships over time, can be categorized as either discrete [7, 8] or continuous [9, 10] based on how they model temporal changes. For instance, in social network analysis, a dynamic graph can represent the evolving patterns of user interactions, where edges appear or disappear as users form new connections or cease communication. Learning the representation of these dynamic graphs is of paramount importance as it encapsulates the interaction and evolution mechanisms of the graph, thereby aiding in understanding and predicting the system behavior.
In contrast to static graph representation learning [11, 12], which focuses primarily on capturing fixed topological features, dynamic graph representation learning faces the dual challenge of preserving evolving topological structures while simultaneously capturing the temporal dynamics of the graph. This dual requirement significantly increases the complexity of model design, necessitating architectures that can effectively capture both spatial dependencies and temporal evolution patterns concurrently. The key challenge lies in maintaining the structural integrity of the graph at each time step while effectively modeling the temporal evolution of both node attributes and edge formations. In this study, we focus primarily on discrete dynamic graph representation learning. This choice is motivated by its widespread applicability across various domains [7, 13, 8] and its efficacy in capturing significant state changes in real-world scenarios, where graph structures often evolve in distinct, observable steps rather than continuous transitions.
A discrete dynamic graph is modeled as a sequence of ordered static snapshots across discrete time intervals, capturing changes in graph topology, including node and edge modifications. Traditional approaches [7, 13, 14] have combined Graph Neural Networks (GNNs) with sequence modeling techniques to capture both structural and temporal information. Recent research has explored more sophisticated architectures, such as transformers and advanced recurrent structures [13], to better handle long-range dependencies and improve inductive learning capabilities. In parallel, continuous dynamic graph modeling methods have gained traction, treating graph evolution as a series of timestamped events. These approaches, exemplified by works like TGAT [15] and CAW [16], leverage attention mechanisms and temporal random walks to capture fine-grained temporal-topological interactions. Advanced techniques such as neural ordinary differential equations [17] and stochastic processes [10, 18] have been employed to model continuous-time dynamics more precisely. Emerging research trends focus on developing methods that can effectively balance the modeling of both short-term and long-term temporal dependencies [7, 19], while also addressing the challenges of scalability and computational efficiency in large-scale dynamic graphs [20]. These advancements aim to provide more comprehensive and nuanced representations of dynamic graph structures, capturing both local and global patterns of evolution.
However, existing approaches exhibit limitations that constrain their effectiveness in dynamic graph representation learning. Firstly, current methods often overlook the dynamic influence of edge temporal states on node features. The temporal evolution of edges—their persistence, emergence, or disappearance—significantly impacts node characteristics, yet many approaches fail to explicitly model these dynamics [21, 22]. Secondly, while recent research [23, 24] has begun to explore transformer architectures to address the over-smoothing problem in GNNs [25, 26], challenges remain in effectively integrating graph structural information with complex temporal dependencies. Moreover, many current approaches struggle to leverage historical node information inductively, limiting their applicability to evolving networks. Consequently, there is a need for an approach that can adaptively integrate temporal and structural information while maintaining generalizability to unseen nodes and edges.
To address the aforementioned challenges, we propose a novel framework for dynamic graph representation learning, named Recurrent Structure-reinforced Graph Transformer (RSGT). At the core of RSGT is a unique graph transformation technique that converts each graph snapshot into a weighted multi-relation graph. This transformation explicitly models edge temporal states through distinct edge types and weights, capturing the variations between consecutive snapshots. Next, we propose a structure-reinforced graph Transformer that introduces a novel structure-aware attention mechanism, modifying the original attention computation to make the Transformer inherently cognizant of graph structures. By operating on our weighted multi-relation difference graph, this enhanced architecture enables the model to simultaneously capture both intricate topological features and temporal dynamics. The structure-aware attention mechanism selectively focuses on salient historical patterns and structural relationships encoded in edge types and weights, effectively learning the complex interplay between graph evolution and temporal dependencies. We then conceptualize dynamic graph representation learning as a sequence of recurrent learning tasks. Starting with initial node features, RSGT continuously updates these representations based on the evolving graph structure and edge states, enabling RSGT to effectively capture long-term dependencies in the dynamic graph. Moreover, the recurrent learning paradigm allows for incremental updates of node representations rather than recomputing from scratch at each time step.
@@
Our extensive empirical evaluations on dynamic link prediction tasks across four real-world datasets demonstrate RSGT’s superior performance in dynamic graph representation learning, highlighting its potential for dynamic graph analysis. In conclusion, the key contributions of this paper can be summarized as follows:
  • 1.
    We propose RSGT, a novel recurrent learning framework for dynamic graphs. By transforming graph snapshots into weighted multi-relational difference graphs, RSGT captures edge temporal states and integrates them with topological structures, enabling a more comprehensive representation of dynamic graph evolution.
  • 2.
    We design a structure-reinforced graph transformer that simultaneously extracts local-global topological features and evolving dynamic characteristics, effectively mitigating the over-smoothing problem while enhancing the model’s ability to capture complex temporal-structural dependencies.
  • 3.
    Extensive experiments on real-world datasets demonstrate RSGT’s superior performance in dynamic link prediction tasks. The results not only validate the effectiveness of our approach in capturing intricate dynamic graph patterns but also highlight its potential for advancing the field of dynamic graph analysis.
The remainder of this paper is organized as follows: Section 2 provides preliminaries and problem formulation. Section 3 details the proposed RSGT framework. Experimental results are presented in Section 4. Related work is reviewed in Section 5, followed by conclusions in Section 6.
2 Problem Formulation
@@
Here, we present the formulations for dynamic graph representation learning through a set of definitions.
@@
A discrete dynamic graph, denoted as $\mathcal{G}$, is an ordered sequence of static graphs represented as $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t}\}$. Here, $\mathcal{G}_{t}=\langle V_{t},E_{t}\rangle$ corresponds to a snapshot of the dynamic graph $\mathcal{G}$ at time slice $t$, where $V_{t}$ is the set of nodes, $E_{t}$ is the set of edges. An edge $e_{ij}^{t}\in E_{t}$ indicates a link between nodes $v_{i},v_{j}\in V_{t}$ at time $t$.
While the dynamic graph captures the overall structure at each time step, it’s crucial to understand how individual edges evolve over time. This leads us to the concept of edge temporal states:
Definition 2 (Edge Temporal States).
The edge temporal states, denoted as $S_{t}$, capture the dynamic nature of edges in a graph over time. For each edge $e_{ij}^{t}\in E_{t}$, its temporal state $s_{ij}^{t}\in S_{t}$ is defined as a tuple $s_{ij}^{t}=\langle tp_{ij}^{t},\omega_{ij}^{t}\rangle$. Here, $tp_{ij}^{t}$ represents the current state of the edge, which can be persistent (the edge exists at both $t-1$ and $t$), emerging (the edge appears at $t$ but did not exist at $t-1$), or disappearing (the edge existed at $t-1$ but no longer exists at $t$). The second component, $\omega_{ij}^{t}$, represents the duration of the current state, indicating the number of consecutive time slices the edge has been in its current state.
Figure 1: Overall framework of the proposed Recurrent Structure-reinforced Graph Transformer (RSGT). (a) Edge Temporal State Modeling of Dynamic Graph: transforming each graph snapshot into a weighted multi-relation graph to model the edge temporal states. (b) Recurrent Temporal Node Feature Extraction: capturing both graph topology and evolving dynamics through a recurrent learning paradigm with the structure-reinforced graph transformer.
While dynamic graphs and edge temporal states provide a framework for describing evolving systems, effectively analyzing these systems requires learning compact representations that capture both structural and temporal information. This leads us to the task of dynamic graph representation learning:
@@
Dynamic graph representation learning aims to learn a mapping function that derives a low-dimensional representation $\textbf{H}^{t}\in\Re^{|V_{t}|\times d}$ ($d\ll|V_{t}|$) by utilizing a succession of snapshots $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t}\}$ and their corresponding edge temporal states $\{S_{1},\dots,S_{t}\}$ up to time slice $t$. The $i$-th row of $\textbf{H}^{t}$, denoted as $\textbf{h}^{t}_{i}$, represents the derived representation of node $v_{i}$ at time slice $t$, encapsulating both structural information and evolutionary dynamics. A typical learning framework can be formulated as:
$\textbf{H}^{t}=g(f(\mathcal{G}_{t},S_{t},\textbf{X}|\Theta_{f}),\textbf{I}_{t-1}|\Theta_{g})$ (1)
where X denotes the initial node features or embedding matrix, $\textbf{I}_{t-1}$ represents the most recent latent temporal state, $f$ is an encoding model for extracting structural and temporal features, and $g$ is a sequence model for discerning evolving patterns. $\Theta_{f}$ and $\Theta_{g}$ are the learnable parameters of $f$ and $g$ respectively.
The effectiveness of node representations learned through this process is crucial for various graph-oriented tasks [27, 10, 28]. One particularly important application is dynamic link prediction, which we use in this study to evaluate our proposed method:
@@
Given a series of snapshots $\{\mathcal{G}_{1},\dots,\mathcal{G}_{t}\}$ up to time slice $t$, the dynamic link prediction task aims to predict whether a candidate edge $e_{ij}^{t+1}$ will exist in the upcoming snapshot $\mathcal{G}_{t+1}$ at time slice $t+1$. This can be formulated as:
$p_{ij}^{t+1}=\tau(\mathcal{G}_{t+1}|\mathcal{G}_{1},\dots,\mathcal{G}_{t},S_{1},\dots,S_{t};\Theta_{\tau})$ (2)
@@
where $p_{ij}^{t+1}$ represents the predicted probability that $e_{ij}\in E_{t+1}$, and $\tau$ indicates the prediction model with parameters $\Theta_{\tau}$.
3 Approach
@@
The overall framework of our proposed RSGT is depicted in Figure 1. It consists of two essential modules: (a) Edge Temporal State Modeling of Dynamic Graph: This module transforms each original graph snapshot into a weighted multi-relation difference graph. It assigns diverse relational types and weights to edges based on their temporal evolution between adjacent snapshots. Specifically, it employs a multi-relational edge type modeling technique to capture short-term dynamics and an edge weight temporal updating mechanism to represent long-term dynamics. (b) Recurrent Temporal Node Feature Extraction: In this module, we introduce a structure-reinforced graph transformer that simultaneously considers the graph topology and the edge temporal states over the weighted multi-relation difference graph. This transformer is used to recurrently update node features, effectively capturing both local structural information and global temporal dependencies. These two modules work in tandem to learn comprehensive and time-aware node representations, enabling RSGT to effectively model the complex dynamics of evolving graphs. In the subsequent sections, we delve deeper into the intricacies of each module.
3.1 Edge Temporal State Modeling of Dynamic Graph
@@
In a dynamic graph $\mathcal{G}$, the feature of a focal node $v_{i}$ can be intuitively perceived as a reflection of its interactions with other nodes $v_{j}\in(N_{i}^{h})^{t}$ that it directly or indirectly engages with [29]. Here, $(N_{i}^{h})^{t}$ signifies the $h$-hop neighborhood of $v_{i}$ at time slice $t$, representing all nodes that can be reached from $v_{i}$ within $h$ steps at time $t$.
@@
Social network theory, as observed in [21, 22], suggests that the strength of relationships between nodes predominantly depends on factors such as historical interaction frequency and duration. This relationship strength, in turn, shapes future interactions, thus playing a significant role in the formation and evolution of the graph. However, capturing these dynamic relationships in a way that effectively informs node representation learning presents several challenges. The interactions between nodes exhibit persistence, emergence, or disappearance across multiple time slices, resulting in perpetual alterations in the temporal states of the edges. Moreover, the temporal state of a specific edge is primarily characterized by two features: short-term dynamics, which represent the edge’s immediate state changes, and long-term dynamics, which constitute the aggregate impact of this edge on the strength of the correlation between its associated nodes. Understanding the evolution of $\mathcal{G}$ thus hinges on accurately learning how distinct edge temporal states impact node features across different time slice snapshots.
To address these challenges and capture these complex dynamics, we propose explicitly modeling edge temporal states, which can provide heuristic information to the downstream feature extraction model, thereby enabling effective learning of node representations. To this end, we propose a heuristic edge temporal state modeling method that transform each $\mathcal{G}_{t}\in\mathcal{G}$ into a weighted multi-relation difference graph $\hat{\mathcal{G}}_{t}$ based on the differences between adjacent graph snapshots. This graph uses different edge types and edge weights to explicitly model the temporal states of edges, capturing both short-term and long-term dynamics. Formally, we define this graph as follows:
Definition 5 (Weighted Multi-relation Difference Graph).
A weighted multi-relation difference graph $\hat{\mathcal{G}}_{t}=\langle V_{t},\hat{E}_{t},S_{t}\rangle$ at time slice $t$ is defined as a graph comprising a set of nodes $V_{t}$, a set of edges $\hat{E}_{t}$ encompassing connections from both current and previous time slices, and a set of edge temporal states $S_{t}=\{\langle tp_{ij}^{t},\omega_{ij}^{t}\rangle\}$, where $tp_{ij}^{t}$ represents the edge type capturing short-term dynamics, and $\omega_{ij}^{t}$ denotes the edge weight reflecting long-term dynamics. Both $tp_{ij}^{t}$ and $\omega_{ij}^{t}$ are modeled based on the differences between adjacent time slice snapshots $\mathcal{G}_{t-1}$ and $\mathcal{G}_{t}$, thereby capturing the temporal dynamics of the graph structure.
@@
where Eq.3 combines edges from consecutive time slices, while Eq.4 categorizes these edges based on their temporal behavior. This categorization allows us to explicitly model the dynamic nature of edges, including those that have disappeared, which is crucial for understanding the evolving graph structure. In $\hat{\mathcal{G}}_{t}$, we provisionally reconstruct the edges that vanished between time slices $t-1$ and $t$. This intervention is based on the premise that the disappearance of an edge does not necessarily signal an immediate termination of inter-influence between the connected nodes. Conversely, the disappearance of an edge might potentially provoke a spectrum of impacts on the two nodes. Consequently, it is crucial for subsequent feature extraction models to explicitly recognize this aspect, as it significantly contributes to comprehending the dynamic nature of the graph.
Secondly, edge weight temporal updating captures the long-term dynamics by assigning and updating weights to edges based on their historical persistence and strength. Specifically, in examining the association between the duration of an edge and the correlation strength of nodes, the edge weight temporal updating procedure assigns a strength coefficient, denoted as $\omega$, to each edge. This coefficient is directly proportional to the duration of the edge, signifying that a longer edge duration corresponds to a higher strength coefficient:
@@
where $\omega^{t}_{ij}$ denotes the strength coefficient of the edge $e_{ij}^{t}$ at time slice $t$, and $k$ is indicative of the edge duration up to the point $t$. The parameters $\alpha$ and $\beta$ control the scaling and growth rate of edge weights respectively. These parameters can be tuned to adjust the model’s sensitivity to edge duration, allowing for flexibility in different application scenarios.
As a result, the dynamic graph $\hat{\mathcal{G}_{t}}$ conserves the original topology while concurrently modeling the temporal state of the edge in a collaborative manner, taking into account the varying edge types and weights. This approach ensures a comprehensive representation of the edge dynamics across different time slices, uniquely capturing both the immediate changes (through edge types) and cumulative effects (through edge weights) of interactions.
3.2 Recurrent Temporal Node Feature Extraction
Figure 2: Structure-reinforced Graph Transformer for temporal node feature extraction. (a) Semantic Encoding Module: capturing global semantic correlations between nodes. (b) Structure Encoding Module: extracting topological dependencies from the weighted multi-relation difference graph, incorporating both original graph structure and temporal information. (c) Temporal Feature Extraction Module: fusing semantic and enhanced topological correlations to generate time-aware node representations.
The Recurrent Temporal Node Feature Extraction process is the key component of our RSGT model, designed to capture both the temporal semantic and structural information of nodes in a dynamic graph setting. This process proposes a structure-reinforced graph transformer (SGT) to iteratively update node representations along the temporal dimension, integrating global semantic correlations, topological dependencies, and edge temporal states across consecutive snapshots. Figure 1 illustrates the process of recurrent temporal node feature extraction over the constructed weighted multi-relation difference graph $\hat{\mathcal{G}}_{t}$. For a sequence of graph snapshots $\{\hat{\mathcal{G}}_{1},\hat{\mathcal{G}}_{2},...,\hat{\mathcal{G}}_{T}\}$, the SGT iteratively updates node representations as follows:
$\hat{\textbf{H}}^{t}=\mathcal{F}_{g_{l}}(\hat{\mathcal{G}}_{t},\textbf{H}^{t-1}),\quad t=1,2,...,T$ (6)